OpenStudy (anonymous):

lim (x^2-9)((x-3)/abs(x-3)) as x approaches 3 (x->3) using squease theorem

OpenStudy (amistre64):

they all squeeze into 0

myininaya (myininaya):

\[\lim_{x \rightarrow 3^+}\frac{(x^2-9)(x-3)}{|x-3|}=\lim_{x \rightarrow 3+}\frac{(x^2-9)(x-3)}{x-3}=\lim_{x \rightarrow 3^+}(x^2-9)=0\]

OpenStudy (amistre64):

split that diff of squares and use it for the other absolute

myininaya (myininaya):

\[\lim_{x \rightarrow 3^-}\frac{(x^2-9)(x-3)}{|x-3|}=\lim_{x \rightarrow 3^-}\frac{(x^2-9)(x-3)}{-(x-3)}=\lim_{x \rightarrow 3-}-(x^2-9)=0\]

OpenStudy (amistre64):

or do we define the abs as (√)^2

myininaya (myininaya):

since the left limit=right limit then \[\lim_{x \rightarrow 3}\frac{(x^2-9)(x-3)}{|x-3|}=0\]